OpenStudy (anonymous):

Find the general solution in terms of a polynomial and a series in powers of x+1: x(2+x)y'' + (2+x)y' + 2y = 0

OpenStudy (anonymous):

After using \[y(x) = \sum_{m=0}^{\infty}a_{m}(x+1)^{n}\] I've found a formula for the constants: \[a_{m+2} = \frac{ a_{m+1}(m+1)+a_{m}(m^{2}+2) }{ (m+2)(m+1) }\] But I'm not sure where to go from here.